OpenStudy (anonymous):

(a^n-2)^3n Im really confused on this whole chapter since I was absent from school.

OpenStudy (anonymous):

Simplify

OpenStudy (bibby):

\[(a^b)^{c} = a^{(b+c)}\] If I remember correctly

OpenStudy (bibby):

it seems powers of a power are multiplied so\[(a^b)^c = a^{b*c}\]

OpenStudy (anonymous):

soo....

OpenStudy (bibby):

a^(n-2)^3n = a^{(n-2)(3n)} idk, why don't you tell me what 3n * (n-2) is

OpenStudy (bibby):

why are you giving up? This is variable multiplication. You learned this a while ago. what is 3x(x-2)? Distribute the outer bit into the parentheses

OpenStudy (anonymous):

3x^2-6x

OpenStudy (anonymous):

It's official OPENSTUDY IS DEAD!

OpenStudy (bibby):

It's official: You're lazy and don't want to apply yourself. you literally just solved the problem...